### v1

# Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment

Abstract
Recent advances in LLMs have greatly improved general-domain NLP tasks. Yet, their adoption in critical domains, such as clinical trial recruitment, remains limited. As trials are designed in natural language and patient data is represented as both structured and unstructured text, the task of matching trials and patients benefits from knowledge aggregation and reasoning abilities of LLMs. Classical approaches are trial-specific and LLMs with their ability to consolidate distributed knowledge hold the potential to build a more general solution. Yet recent applications of LLM-assisted methods rely on proprietary models and weak evaluation benchmarks. In this survey, we are the first to analyze the task of trial-patient matching and contextualize emerging LLM-based approaches in clinical trial recruitment. We critically examine existing benchmarks, approaches and evaluation frameworks, the challenges to adopting LLM technologies in clinical research and exciting future directions.
Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment
Shrestha Ghosh1   Moritz Schneider2   Carina Reinicke2   Carsten Eickhoff1 1University of Tübingen, Germany 2Boehringer Ingelheim, Germany 1{first.last}@uni-tuebingen.de   2{first.last}@boehringer-ingelheim.com
1 Introduction
Clinical trials evaluate the effects of an intervention on human health. Selecting the precise and required size of patient population is crucial for trial completion. According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time Clinical Trials Arena (2012); Williams et al. (2015); Pharmaceutical Technology (2019). While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022), thereby hampering the rate at which potentially life-saving therapies or new vaccines can be introduced.
Figure 1: Steps in a typical patient recruitment process vs. our proposed flow with expert interaction.
Clinical trials have a dual nature, consisting of both trial-agnostic and trial-specific requirements, making it challenging to design generalized approaches. This has traditionally resulted in highly specialized matching processes that work for a particular trial or patient population. The resulting patient recruitment process is linear, as illustrated (in orange) in Figure 1, leaving little room for expert intervention and feedback. It starts with extracting information from a trial description and available patient records, then matching explicit trial eligibility criteria and patient data to determine patient eligibility. Finally, experts review the selected patients before contacting them.
An expert reviews several hundred patients Ni et al. (2015) and can end up spending hours on one patient, incurring significant costs Penberthy et al. (2012). In their study on five trials, Penberthy et al. (2010) showed that automated matching, via table queries and string-matching algorithms, saved between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation. As more patient data becomes digitally available, research into applying machine intelligence and developing explainable systems for clinical researchers to correctly interpret model predictions has gained traction Murdoch and Detsky (2013); Payrovnaziri et al. (2020); von Itzstein et al. (2021).
Generative LLMs can serve as knowledge aggregators and streamline clinical trial design and trial-patient matching (in purple in Figure 1). However, the clinical trial recruitment process has seen relatively low adoption of transformer language models He et al. (2020); Harrer (2023); Lu et al. (2024). Instruction fine-tuned LLMs for generating explainable trial-patient matches has received attention only since the past one year Jin et al. (2024); Nievas et al. (2024); Rybinski et al. (2024); Wornow et al. (2025). Given the sensitive nature of clinical trials and the rapidly evolving landscape of LLM technology, there is no prior work surveying this area. We fill this gap by critically examining the trial-patient matching approaches, the method used, the benchmarks available and the evaluation strategies. The existing systematic reviews of NLP methods and their performance Kim and Quintana (2022); Idnay et al. (2022); Kantor and Morzy (2024) are desiged to follow strict inclusion/exclusion criteria and search strategies Khan et al. (2003); Moher et al. (2009). This practice, essential in medical research for measuring the association between intervention and outcomes, misses the larger picture of contextualizing LLMs in clinical trial recruitment settings. In contrast, this work follows the perspective of surveys in computer science (see Appendix B) to provide a more comprehensive view of the role of LLMs in clinical trial recruitment.
Our Contribution
Cohort discovery has been a long-standing problem and LLM-assisted methods to improve trial-patient matching has only recently gained momentum. In this work, we formalize the problem of trial-patient matching, and differentiate between two major approaches (trial-centric and patient-centric), the associated tasks (classification vs. ranking) and their effect on the choice of benchmarks (longitudinal vs. short patient descriptions, single vs. multi-trial). We identify the important bottlenecks in data availability and task formalization. We introduce a taxonomy of errors in LLM-generations for trial-patient matching. Finally, we provide actionable steps for the community on four promising directions.
2 Background
Clinical Trial Recruitment
Also known as patient recruitment/enrollment/(pre-)screening, it is the process of finding relevant patients (or cohort) for a clinical trial. The steps in patient recruitment, as shown in Figure 1, include data sourcing, information extraction, matching, and expert review.
Eligibility Criteria
They are explicit conditions for (inclusion) or against (exclusion) enrollment, usually expressed in natural language and designed by trial investigators to select cohorts. These conditions have varying difficulty levels, ranging from simple tabular operations (age or gender requirements) to semantic matching (medical condition or medications) and inference (reasoning with negation, patient history).
Patient Records
They refer to any form of data, in a combination of structured and unstructured formats, used to describe a patient. When digitized, it is often called an Electronic Health Record (EHR). This data includes the patient’s demographic information (age, gender, ethnicity), clinical notes, medications, billing history, and test results, that are collected over a period of years.
Biomedical NLP and LLMs
Transformer language models have greatly improved information retrieval and extraction tasks. There exist specialized resources, such as, PubmedQA Jin et al. (2019), MIMIC-III & IV Johnson et al. (2016, 2020) and BioRED Luo et al. (2022), to train encoder language models Radford (2018); Devlin et al. (2019); Raffel et al. (2020) on medical domain (e.g., BioBERT Lee et al. (2020), PubmedBERT Gu et al. (2021), Criteria2Query2.0 Fang et al. (2022), MedCPT Jin et al. (2023)). Yet, Dorfner et al. (2024) show that fine-tuning generative LLMs on biomedical domain has limited relative performance gain. In fact, Alber et al. (2025) show that medical misinformation encountered during the fine-tuning affect downstream tasks. The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, observed a shift from domain-specific small models towards larger general-purpose LLMs Lu et al. (2024). Even among LLMs, the OpenAI Achiam et al. (2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced LLaMA Touvron et al. (2023) or Mistral Jiang et al. (2023) models. Finally, Bedi et al. (2025) report that of more than 519 biomedical studies using LLMs, less than 5% used real patient data, tasks like summarization are understudied and fairness, bias, uncertainty and deployment considerations are rarely measured.
State of Adoption of NLP Advancements
Historically, patient recruitment has been a manual process with minimal automation via pre-defined database queries Thadani et al. (2009), and, more recently, automatic translations of free text into database queries Yuan et al. (2019); Fang et al. (2022); Park et al. (2024). A recent review reported low adoption rates of AI, particularly NLP, in clinical trials Kantor and Morzy (2024). Another review on understanding the role of NLP systems in patient recruitment found just 11 studies that evaluated NLP systems in clinical settings Idnay et al. (2022). Heterogeneous outcomes, diverse results and a dependence on small retrospective data drive the gap in NLP research and their adoption in real-world settings Idnay et al. (2024). A study by Idnay et al. (2023), investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies. Interestingly, when Corbaux et al. (2024) categorize trial matching tools for oncological trials as manual, semi-automatic and automatic, automatic methods still fall in the research and development phase, yet to be commercially available.
Review Methodology
We searched for relavant papers since 2019, using the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt”. We included method and dataset papers using LLMs and excluded papers that used LLMs in biomedical NLP (broader scope) or other health care applications (not clinical trials). Starting with the citations in the LLM papers we tracked the pre-LLM literature by recursively tracing their citations in newer publications. Since the topic “clinical trial recruitment” is vast, this approach efficiently captured the evolution of the predictive methods in this field.
3 Trial-Patient Matching
Trial-patient matching assigns a trial-patient pair one of the three labels of eligible, ineligible and irrelevant based on the patient record and eligibility criteria. Importantly, a trial-centric approach, taken by a trial investigator, matches patients from a hospital database to a specific trial. Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a patient.
Given sets of inclusion and exclusion criteria ($C_{inc},C_{exc}$) from a trial and a set of patient data, $P$, we formalize the trial-patient matching problem, $M$, to predict one of the labels Ineligible (Inel.), Irrelevant (Irr.) or Eligible (Eli.) by aggregating criterion-level binary matches $M^{\prime}(c,P)$.
$\!M(C_{inc},C_{exc},P)\!=\!\left\{\!\begin{array}[]{ll}\!\textit{Inel.},\!&\!\exists c\in C_{exc},M^{\prime}(c,P)\\ \!\textit{Irr.},\!&\!\exists c\in C_{inc},\neg M^{\prime}(c,P)\\ \!\textit{Eli.},\!&\!\neg(\textit{Inel.}\lor\textit{Irr.})\end{array}\!\right.$
This induces a priority, such that, a patient satisfying any exclusion criteria becomes ineligible, regardless of inclusion criteria matches. In Section 6, we elaborate on the lack of a formal task definition and its effects on trial-centric matching.
Task Benchmark Size Eval. Metrics Best Score Cohort
selection
2018 N2C2
Stubbs et al. (2019)
#patients = 288
#records per patient = 2-5
#tokens/patient = 2711
#trials = 1
#criteria = 13
Micro-averaged:
P, R, F1
0.91 micro F1
Trial
ranking
Koopman and Zuccon (2016) #patient summaries (22 words avg) = 60
#patient description (78 words avg) = 60
#keywords (4.4 words avg) = 489
#trials = 204,855
#relevant matches = 685
#relevance judgements = 4000
MRR,
P@5,
adaptive precision
0.3 MRR
$<$ 0.2 P@5
TREC CT Track 2021
Soboroff (2021)
#patient descriptions = 75
#trials = 375,580
#relevant matches = 5,570
#relevance judgements = 35,832
NDCG@10,
P@10,
MRR,
R-Precision
0.84 NDCG@10
0.748 P@10
1.0 MRR
TREC CT Track 2022
Roberts et al. (2022)
#patient descriptions = 50
trials = 375,580
#relevant matches = 3,949
#relevance judgements = 35,394
NDCG@10,
P@10,
MRR,
R-Precision
0.61 NDCG@10
0.50 P@10
0.32 R-Precision
0.72 MRR
TREC CT Track 2023
Rybinski et al. (2024)∗∗
#patient tables = 40
#disease templates = 8
#trials = 451,538
#relevant matches = 11,963
#relevance judgements = 34,931
NDCG@10,
P@10,
MRR
0.81 NDCG@10
0.73 P@10
0.78 MRR
Best scores are aggregated from the Appendix of the proceedings of TREC 2021. ∗∗Borrowed from Rybinski et al. (2024) as TREC CT 2023 does not have a published track overview.
Table 1: Overview of public evaluation benchmarks for trial-patient matches covered in this survey.
4 Benchmarks and Corpora
We analyze five public benchmarks in context of the corpora and their evaluation metrics.
  • 2018 N2C2 Cohort Selection for Clinical Trials Stubbs et al. (2019). [^1] Currently unavailable as of 2024 Nov 6.
  • Trial ranking Koopman and Zuccon (2016).
  • Text REtrieval Conference Clinical Trial (TREC CT) [^2] Available at https://www.trec-cds.org/. tracks 2021, 2022, 2023 for ranking trials.
Table 1 provides an overview of the benchmarks.
4.1 Trial Data
ClinicalTrials.gov is one of the largest online database of clinical studies, including clinical trials and other observational studies. As of November 2024, it lists more than 390,000 clinical trials, thousands of which are active. These studies are submitted by investigators from over 200 countries. The euclinicaltrials.eu is another online registry for trials conducted in the European Union. It comprises more than 50,000 trials, with more than 7000 being active. Every trial comprises a study description, eligibility criteria and study plan among other details. The ClinicalTrials.gov database forms the primary source of trials in many academic studies.
As we will see in Section 5, earlier trial-specific approaches manually separated discrete data from free-text search. However, this separation is difficult to scale, and newer approaches focus on semantic matching. The N2C2 cohort selection benchmark is based on a single trial. Hence, systems specifically model the trial criteria. The TREC CT benchmarks source the ClinicalTrials.gov database for the task of trial ranking. However, additional annotation of the eligibility criteria denoting their type or complexity is unavailable.
4.2 Patient Data
Since patient data contains sensitive information, additional measures are required to remove personal information in order to release them as public repositories. The MIMIC database Johnson et al. (2016, 2020) is the largest public database of patient information. The latest version, MIMIC-IV, contains rich data on hospital admissions, medications, demographic data and notes of 299,712 de-identified patients. This dataset is not used in any of the five benchmarks. One reason might be a significant challenge in obtaining confirmed trial-patient match labels, which is easily available retrospectively from existing hospital records. The N2C2 cohort selection benchmark Stubbs et al. (2019), contains 288 de-identified longitudinal records of patients. The TREC CT benchmarks use synthetic patient profiles, comprising short textual patient descriptions and keywords defining cohorts.
4.3 Evaluation Metrics
Standard IR metrics (precision, recall, F1) and their derivatives (precision at rank (P@k), normalized discounted cumulative gains (NDCG@k), R-precision, mean reciprocal rank (MRR) and adaptive precision Moffat et al. (2015)) are the most popular evaluation metrics. Accuracy, specificity, and workload are also prevelant in the medical community. Notably, researchers in the medical domain have stressed the lack of definition and consistency of ground truth data to evaluate NLP tools in clinical trial matching Kim and Quintana (2022). The trial-patient ground-truth labels are obtained manually either from medical experts Koopman and Zuccon (2016); Soboroff (2021); Roberts et al. (2022) or retrospectively from past trials Stubbs et al. (2019). The TREC CT annotations were created from the top-k results pooled among all participating teams, resulting in tens of thousands of relevance judgments. Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient Roberts et al. (2022). In the 2023 edition, the relevance judgments were more balanced between the three labels Rybinski et al. (2024). Measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent in the current works.
5 Trial-Patient Matching Methods
We broadly categorize systems into classical and LLM-based approaches and further group them by their direction of approach, trial/patient-centric. The LLM-based approaches focus on tasks, such as, criterion/trial-level predictions and relevance ranking, which in principle is direction-agnostic. Appendix A lists the works cited in this paper.
5.1 Classical Approaches
Trial-centric Matches
Majority of classical models use this approach build trial or cohort specific models. They follow a two-step method, starting with manually-designed rule-based discrete data filters, followed by sub-word matching Penberthy et al. (2010), vector matching Ni et al. (2015) or machine-learned classifiers Tun et al. (2023) for matching eligibility criteria or cohort keywords. Ni et al. (2015) showed a retrospective workload reduction to 50 patients down from 163 per trial. IBM Watson for clinical trial matching (IBM-CTM) is a proprietary commercial tool for trial-patient matching, which uses neural NLP techniques to match eligibility criteria to patients Helgeson et al. (2018). While one study on 3 trials measured a reduction in workload from 110 minutes to 86 minutes for screening 90 patients Beck et al. (2017), another study on 10 clinical trials and 102 patients evaluated the agreement in matches between IBM-CTM and expert annotation to be above 80% Alexander et al. (2020). Another line of work that focuses on defining a cohort by translating natural language criteria into formal query language for instantly querying patient databases Yuan et al. (2019).
Patient-centric Matches
Ni et al. (2015) create bag-of-words vectors to retrieve top trial matches for a patient and show a retrospective workload reduction (4 vs. 42 trials per patient) and 33% precision on a subset of 127 patients. The workload increases to 20 trials per patient and the precision drops to less than 4% when generalized to 215 patients. Zhang and Demner-Fushman (2017) train an SVM classifier on the top unigram and bigram features to predict if a trial is eligible for a cohort (e.g., HIV patients). Additional features derived from NER and lemmatization yielded small gains in performance.
5.2 LLM-based Approaches
These approaches do not rely on trial-specific heuristics. Yet, as systems are evaluated on a few disease-specific trials, the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al. (2023), the generalizability of LLMs across diseases is unclear. Moreover, when compared, LLMs are less effective in making trial-level predictions (low F1 scores) than criterion-level ones.
Trial-Level Prediction
Wong et al. (2023) used GPT4 to structure eligibility criteria into logical expressions. Despite lagging behind expert systems, which use heavy manual intervention, they demonstrate the readiness of general-purpose LLMs for cohort-specific NER (oncology in this case) and logical structuring. Zero-shot and few-shot LLM prompts outperformed the state-of-the-art biomedical NER, SciSpaCy Neumann et al. (2019).
Yuan et al. (2023) prompt LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility. The trial-level prediction is then composed by asserting that all inclusion criteria and none of the exclusion criteria match. On six trials for stoke patients, the model achieved 0.91 F1 at criterion-level vs. 0.81 F1 at trial-level.
Criterion-Level Prediction
Hamer et al. (2023) use OpenAI’s text-davince-003 model to first filter criteria that applies to a patient, then justify whether applicable criteria are met and to finally predict criterion-level labels based on previous explanations. This effectively reduces the total number of eligibility criteria (down to 300 from originally 4,100) a physician has to assess before recommending trials to a patient.
Wornow et al. (2025) prompt LLMs with trial-patient pairs to generate structured JSON strings of criterion-level predictions, explanations with evidence and a self-declared confidence. The content of the prompt ranges from the entire patient record and a trial to a single criterion and relevant record snippets. GPT4 achieves a micro-F1 of 0.93 vs. 0.91 points by the prior state-of-the-art on the 2018 N2C2 cohort dataset. The RECTIFIER Unlu et al. (2024) employs a RAG-based approach to find top-k relevant patient data stored as vector embeddings. The criteria and relevant records are prompted to the LLM to generate either Yes or No. Evaluated on just a single trial and a pool of 3000 private patient records, with clinical notes spanning three years, RECTIFIER achieved an accuracy higher than trained non-licensed staff overall (92.7% vs. 89.1%), especially for the Symptomatic Heart Failure criterion (97.9% vs. 91.7%).
Trial Ranking
One line of work investigates the role of neural re-rankers for ranking relevant trials. Pradeep et al. (2022) generate multiple independent queries from a patient description to retrieve relevant trials. The trial and patient description is then passed to a fine-tuned T5 model to generate a relevance label. Zhuang et al. (2023) propose a multi-stage pipeline that involves a hybrid sparse-dense retriever, initialized with PubmedBERT-base and fine-tuned to retrieve the top 1000 trials. A cross-encoder, initialized with PubmedBERT-large, is trained to re-rank these results. Finally, the top-20 trials are re-ranked using GPT4. Rybinski et al. (2024) leverage LLMs for query reformulation and trial re-ranking to outperform traditional BM25 and BERT-based retrieval models. They show that BM25 is an effective initial retriever, surpassed by BERT-based neural re-rankers.
Jin et al. (2024) use GPT4 to first extract query keywords to retrieve top-k trials and then prompt GPT4 with patient data and eligibility criteria to generate criterion-level labels and explanations. They use a combination of manually-defined and GPT4 prompted scores to determine a trial-level score for final raning. Their TrialGPT performs better on the TREC datasets compared to smaller domain-specific transformer models and show a screening time reduction by 42%. Trial-LLama by Nievas et al. (2024) extend this study by extensive fine-tuning of open-source LLaMA models.
6 Discussion
Annotated Corpora
There have been few attempts to analyze and structure the clinical trials repository. As a result, finding similar trials and analyzing systematic failure cases is notoriously difficult. Chia, a manually annotated corpus of 1000 clinical trials created by Kury et al. (2020) has criterion-level annotations of entities, relations and Boolean logic graphs. Such annotations on trial-patient matching benchmarks can help us measure model fairness. Scaling patient data collections is also challenging as evident from Table 1, where patient records are in the hundreds. Individual records, curated manually, range from keywords via short patient summaries to documents over a time period. The format ranges from free text via semi-structured in-line tables to full-fledged databases. The rich and large MIMIC dataset, remains largely untapped for the task of trial-patient matching. LLMs provide a promising avenue for solving the automatic annotation at scale.
Grouth-truth Labels
Access to trial-patient eligibility labels, i.e., {eligible, ineligible, not-relevant}, is integral for training and evaluation. As they are obtained from human experts, they are difficult to scale and take time. Hence, many systems are evaluated retrospectively on patient data from private repositories such as clinics that maintain participation records in past trials Wong et al. (2023); Yuan et al. (2023); Hamer et al. (2023); Unlu et al. (2024). The largest public annotations from the TREC CT tracks have an average of 700 trial annotations per patient description. Bedi et al. (2025) found that of the 519 studies between 2022 and 2024 that evaluated LLM applications in healthcare, only 5% used real patient data. Jointly, the major public data sources, e.g., ClinicalTrials.gov and MIMIC datasets, present an exciting opportunity to develop generalizable matching methods.
Formalization
We formalize the task of trial-patient matching in Equation 3 through simple set operations. Surprisingly, except Yuan et al. (2023), none of the related work explicitly formalize the task. This coincides with a lack of aggregation strategies at the criterion-level Hamer et al. (2023); Wornow et al. (2025); Unlu et al. (2024). Formalizing the task, guides Yuan et al. (2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria. The criterion-level labels by Yuan et al. (2023), while not being directly interpretable, are aggregated as in Equation 3 for the trial-level labels. Wong et al. (2023) could potentially explain prediction via logical component mismatches, though the current F1 measures are below 50 points. Clinical experts, who analyze matches for the final cohort selection, strictly follow Equation 3. Criterion-level predictions and explanations have shown to reduce screening time by 42% Jin et al. (2024).
Core Limitations of LLMs
According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood. In Section 7, we discuss how knowledge recall and lack of self-assessment surface through reasoning errors. These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
Societal Impact
Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al. (2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al. (2024). There is a lack of such analyses and mitigation methods in trial-patient matching and a lack of diverse datasets. The popularity of propreitary models for biomedical tasks Lu et al. (2024) including trial-patient matching Wornow et al. (2025); Jin et al. (2024) further ethical concerns on handling sensitive patient data. Alber et al. (2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation. More research is required to determine the feasibility of locally deploying open-source models, the associated costs of maintaining them and keeping them up-to-date.
Figure 2: Taxonomy of errors in LLM-generations.
7 Taxonomy of LLM Errors
Since there is no one accepted taxonomy of errors, we come across semantically overlapping error types that are used to evaluate LLM generated explanations. For instance, incorrect reasoning and lack of knowledge are recognized as independent error types in Jin et al. (2024); Hamer et al. (2023); Nievas et al. (2024), even though the latter often leads to the former. In the proposal by Liévin et al. (2024), reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge. We therefore identify where errors occur and break them down further by an identifiable source of the error (Figure 2).
Information Isolation Errors
These are errors in information extraction from patient data or trial criteria falls. This includes missed or incorrect NER, measurements (numeric or unit), temporal scopes and negations. LLMs are good at recognizing entities and measurements, but still struggle with negations Nievas et al. (2024).
Reasoning Errors
Any error in the explanations provided by LLMs fall into this category. The most common source is insufficient explicit data. Basically, LLMs fail to draw logical conclusions from given data, while an expert can. This stems from previously unseen data (“lack of knowledge”) or the inability to recall prior information (“lack of implicit reasoning”) or the inability to infer from context (“insufficient data”).
The second source of reasoning errors is lack of self-assessment. The LLM contradicts explicit information in the prompt. The error occurs when knowledge is wrongly recalled or knowledge is correctly recalled, but contextual information is not applied, resulting in wrong reasoning. This is often referred to as “incorrect knowledge”.
The third source are premature decisions made by an LLM when the criteria actually require information that is missing, such as, expert opinion.
Inconsistency Errors
Generating explanations to arrive at final answers unlock LLM’s reasoning capacities Wei et al. (2022). Even so, the explanation and the prediction can be inconsistent. LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in Nievas et al. (2024)). The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to report unless explanations for correct predictions are evaluated as well. It may seem as if the LLM is taking a shortcut, hinting towards a bias that the model picked up during training. Both these cases negatively impact transparency and accountability of the matching system.
8 Future Directions
Clinical trial recruitment spans multiple interactive parts. As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al. (2024); Qiu et al. (2024), we focus on four promising directions.
Trial Search
Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses. This task goes beyond keyword search to match eligibility criteria of two trials. This problem has seen little activity since clustering using lexical features Hao et al. (2014). Analogous systems, such as SemEHR Wu et al. (2018), exist for interactive semantic search over patient EHRs. Sun et al. (2022) combine multiple similarity measures to account for categorical features (study type, phase, and so on) and lexical features (condition name, intervention name, outcome measures, eligibility criteria) to find similar trials. Jullien et al. (2023) deploy LLMs to find trials that match short descriptions using textual entailment and ranked evidences. However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
Interactive Trial Design
Investigators design trials independent of patient data. This creates a semantic gap between eligibility criteria and patient data. LLM agents have the potential of bridging this gap by suggesting data models underlying patient data for structuring eligibility criteria early on. This collaborative idea is not new Luo et al. (2013), however, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al. (2017); Sun and Loparo (2019); Liu et al. (2021a); Dasgupta et al. (2020). Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria Clinical Trials Arena (2022). LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate Liu et al. (2021b).
Collaborative Trial Planning
Eligibility criteria have a range of difficulty levels. A preemptive task can interact with patient records to estimate the size of the patient pool on the most restrictive criteria, say the presence of a rare genetic disorder. As patient records progressively follow uniform data models, it is more efficient to separate and structure criteria that require tabular operations (e.g., via structured queries) from criteria that require reasoning (temporal, numerical, negation), and to separate criteria that require expert analysis from being wrongly used in semantic searches Mozannar and Sontag (2020). Markey et al. (2025) showed promising results on content relevance and suitability with room for improvement in logical reasoning and provenance when using LLMs to generate trial protocols. Training LLM agents to identify and distribute tasks and then aggregating is a fertile ground for innovation.
Explainable Matches
In the context of LLMs, accurately explaining black-box predictions in human-understandable form becomes very challenging Zhao et al. (2024a). Explainability research in clinical trial matches has been limited to prompting LLMs and manually evaluating chain-of-thought generations post-hoc, which is only one of the many facets of evaluating explainability Nauta et al. (2023); Bodria et al. (2023); Chen and Eickhoff (2024). Moreover, it is unclear how helpful these explanations are, especially, as discussed in Section 7, since they give rise to inconsistency errors and are categorized into ambiguous error labels.
9 Conclusion
The ability of generative LLMs to generalize comes with a trade-off between reasoning and consistency. This work is the first to survey the rapidly evolving role of generative LLMs in clinical trial recruitment. We provide a modern perspective on the challenges in adopting LLMs to clinical research and point out promising directions for future work.
10 Limitations
This survey focuses on the role of advances in NLP in the critical domain of clinical trial recruitment. Given that this is a rapidly evolving field, we have made our best effort to include a comprehensive view of available resources and methods. It is possible that more sophisticated methods using the latest technology already exist (e.g., in the form of proprietary products), but are not yet made public.
Appendix A Trial-Patient Matching Systems
We show all the classical systems in Table 2 and the LLM-based systems in Table LABEL:tab:systems_overview_llm, by their task, method, data source characteristics and limitations.
Appendix B Systematic Reviews versus Surveys
A systematic review is a common practice in the medical research community with standardized reporting guidelines. The Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) Vrabel (2015) is the most widely used guideline. Just as finding patients given a trial eligibility criteria, this guideline requires strict inclusion and exclusion criteria to be applied to scientific literature to be included in a systematic review.
In contrast, the guidelines for writing surveys in our community, for instance, as outlined for authors by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
ACM Computing Surveys (long paper)
A paper that summarizes and organizes recent research results in a novel way that integrates and adds understanding to work in the field. A survey article assumes a general knowledge of the area; it emphasizes the classification of the existing literature, developing a perspective on the area, and evaluating trends.
TACL (excerpt)
They should thus not simply be a descriptive enumeration of the contents of papers, but draw broad themes and (importantly) provide new insights on the topic. These insights should be major contributions of the submission.
ARR (note)
all papers are expected to include reviews of related literature. This category is meant for the papers that go beyond that, e.g. in scope or in establishing new interdisciplinary connections.
Direction Method Patient Source Trial Source Limitations Trial-centric Penberthy et al. (2010) Discrete-data filter and sub-word matching Availability:
Private
Size 282-2,112
Mode: text, tables
Additional annotation: expert review
Availability:
public (1); unknown (4)
Size: 5
Type: Specific
Manually coded eligibility criteria.
String similarity for keyword search.
Yuan et al. (2019) (Criteria2Query)
NL to structured queries
Information extraction, normalization
Mapping to pre-defined cohort templates
No patient data Availability:
Public
Size: 10
Type: Varying
Additional annotation: 125 criteria sentence, 215 entities, 34 relations, 137 negations, 20 attributes
Disease-specific NER.
Query templates cannot deal with missing attributes.
Entity normalization into a small set of 2000 concepts.
Tun et al. (2023)
Rule-based filters
Similarity scores per criteria as features to a classifier
Availability:
Private
Size: 40,000
Mode: text, tables
Additional annotation: 109 patients labels
Availability:
Unknown
Size: 1
Type: Specific
Limited to a single trial
Rules and classifiers are not robust to changes in criteria
Ni et al. (2015)
Discrete-data filter
Index trial and patient bag-of-words vectors
Return top vector matches for a trial
Availability:
Private
Size: 215
Mode: text, tables
Additional annotation: historical match, expert review
Availability:
Public
Size: 55
Type: Specific
12.6% precision
Patient-centric Ni et al. (2015)
Return top vector matches for a patient
(same as previous row) 4% precision
Zhang and Demner-Fushman (2017)
Bag-of-words feature vector
SVM classifier
No patient data Availability:
Public
Size: 2461
Type: 891 Specific; 1570 Varying
Additional annotation: Trial labels
Cohort-specific model
No real patient data considered
Closest to keyword search (trials to cohort)
Public source for clinical trials: https://clinicalTrials.gov
Table 2: Overview of classical systems covered in this survey: direction, methods, data sources and limitations.
Task Method Patient Source Trial Source Limitations
Trial-level prediction Wong et al. (2023)
LLM prompted with a template to generate structured forms of eligibility criteria
Availability: Private
Size: Unknown
Mode: Structured
Additional Annotation: 523 trial-patient historical labels; 68,485 trial-patient new labels
Availability: Public∗∗
Size: 53
Type: Specific
Clinical Trial eligibility limited to the first 40 lines.
Trial-level performance is very low (F1 score range:[29.6-48])
Yuan et al. (2023) Prompt LLMs to reformulate eligibility criteria Availability: Private
Size: 825
Mode: Longitudinal text records
Availability: Public∗∗
Size: 6
Type: Specific
Additional Annotation: 100,000 criterion-patient labels
Criterion-level performance metrics 10 points higher than trial-patient.
High variance between different trials: F1 range: 0.39 - 0.96.
Criterion-level prediction Hamer et al. (2023)
1-shot prompt to LLM for criteria level prediction with explanation
Availability: Private (synthetic patient profile)
Size: 10
Mode: Short text
Type: Specific
Availability: Mixed (clinicalTrials.gov, EudraCT)
Size: 146 clinical trials
Type: Specific
Additional Annotation: Expert review
Majority of criterion-level errors due to incorrect reasoning (91%)
Unlu et al. (2024)
RAG-based
Availability: Private
Size: 2,276
Mode: Tables; text
Availability: Public∗∗
Size: 1
Type: Specific
RAG pipeline not published. Evaluated on a single trial.
Wornow et al. (2025)
Vector database of patient notes
Prompt (criteria, top-k patient notes) pair to LLM
Availability: Public (2018 N2C2: Cohort Selection) Evaluated on one trial
Doesn’t consider exclusion criteria
67% incorrect decision despite correct reasoning.
Trial ranking Pradeep et al. (2022)
Synthesis queries for initial trial retrieval.
Fine-tune T5 to generate relevance label based on patient description and trial data
Availability: Public (TREC CT 2021) Zero-shot relevance ranking is only slightly better than BM25. Fine-tuned models need re-training on new data or format.
Zhuang et al. (2023)
Hybrid sparse-dense retriever for top-1000.
Cross-encoder re-ranker
GPT4 for top-20 re-rank
Availability: Public (TREC CT 2023) High NDCG, 0.51 precision, 0.38 recall (very low)
Validation on textual patient note does not transfer well to testing on structured patient note
Rybinski et al. (2024)
Multi-stage retriever using LLMs
Availability: Public (TREC CT 2021, 2022, 2023) High latency of GPT4 vs a small performance boost
Jin et al. (2024)
Hybrid trial filtering with BM25 and MedCPT Jin et al. (2023)
Criterion-level LLM prediction aggregated to trial scores
Availability: Public (Koopman and Zuccon (2016), TREC CT 2021 & 2022)
Size: 58; 75; 50
Mode: Short text
Additional annotations: 3,621; 26,149; 26,581 trial-patient matches
Availability: Public∗∗
Size: 375,580
Additional annotation: 1,015 criterion-patient labels
User study sample size is small
Does not handle longitudinal patient data
LLM aggregation requires LLM to perform mathematical reasoning and computations
Nievas et al. (2024)
Extends Jin et al. (2024) to open-source LLMs
(same as Jin et al. (2024))
Additional annotation: Patient sentence supporting eligibility label Additional annotation: 500 criteria labels on: eligibility and difficulty.
Significantly high fine-tuning costs
EudraCT (European Union Drug Regulating Authorities Clinical Trials) is the European clinical trials database EudraCT .
∗∗ Public source for clinical trials: clinicalTrials.gov.
Table 3: Overview of LLM-based systems covered in this survey: tasks, methods, data sources used and limitations.